OpenStudy (anonymous):

NEED HELP ASAP! 18. Simplify the expression: -x(6x - 7) A. -6x^2 + 7x B. 5x^2 - 6x C. 6x + 7x D. -6x - 7x

OpenStudy (tkhunny):

Well, do it. What's your plan. Never post a problem with no work shown.

OpenStudy (anonymous):

I am here since I need help on the problem. That's why I came here, to learn how to solve the problem.

OpenStudy (zale101):

Just post any work you have done regarding this question. it doesn't matter if it's wrong. We can help you out and to fix the problems where you made mistakes on:)

OpenStudy (tkhunny):

You MUST have SOME idea. Give us ANYTHING. We'll be delighted.

OpenStudy (anonymous):

I think the answer is C since -x * -6x would be 6x then -x * -7 would give you a positive so 6x + 7x?

OpenStudy (zale101):

-x * -6x But you had just positive 6x \(-x(\color{red}6x - 7)\)

OpenStudy (anonymous):

So -6 so either A or D?

OpenStudy (tkhunny):

a(b+c) Use Distributive Property of Multiplication over Addition a(b+c) = a*b + a*c If you get 6x, which is where you started, something went wrong.

OpenStudy (anonymous):

So A?

OpenStudy (tkhunny):

Let the notation help you. Go ahead and write things out. -x(6x - 7) = (-x)(6x) + (-x)(-7) = -6x^2 + 7x You can demonstrate the entire process in just a few characters.

OpenStudy (anonymous):

Yay. Thank you so much.

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

1. Which polynomial does the model represent? A. -x^3 + 2x^2 - 3x + 2x B. x^2 + x - 2 C. -x^2 - x + 2 D. -x^2 + x - 2

OpenStudy (tkhunny):

One question per thread, please. As stated in my original remarks on this thread, do not ever post a problem with no work shown. Stop making us work to get something out of you. Offer it freely.